# Witnessing Disorder in Quantum Magnets

### v2

@@ VIII.5 What can be useful experimentally?
For 1d critical states, the low-temperature behavior of concurrence can access central charge information. In particular, the nearest neighbor concurrence of an $SU(2)$ symmetric ground state can be related to the internal energy, and this allowed us to extract the central charge of the TLL state. On the other hand, for RS state, while the order of disorder-average effects how concurrence is distributed among the sites, the universal information about the state should still be accessible if it is contained within the 2-partite entanglement. This universal information is the effective central charge. Since experimentally, the concurrence distribution for the RS state can be nearest-neighbored just as the TLL, we thought that perhaps one might be able to access the effective central charge for the RS state. This is based on the assumption that there exists a form of generalized c-theorem for the RS state[92]. This is what we put to test in Fig. 6 and see some positive indications. However, we admit that this is conjectural at best.
@@
On the other hand, multi-partite entanglement witnesses like the QFI or the equal time structure factor are invaluable markers for not only identifying the existence of multi-partite entanglement but also how disorder affects their growth. We saw this by the marked difference in the low-temperature behavior of the equal time structure factor in the TLL and RS state. This behavior can be experimentally useful in identifying the role of disorder in spin chains.
VIII.6 Limitations to detection of multi-partite entanglement
@@
Measures like QFI, when accessed via the dynamic structure factor, can only detect multi-partite entanglement if the underlying state shows correlations peaked around specific $\mathbf{q}$ points. Due to the sum rule, this structure constraints the visibility of multi-partite entanglement detection at other $\mathbf{q}$ points. However, such peaked structure factors are not universal. For instance, in Kitaev and Kagome QSLs, there are no special $\mathbf{q}$ points with sharp features [99, 100, 101]. Therefore, multi-partite entanglement cannot be accessed using INS in these cases. This does not imply the absence of multi-partite entanglement in these states, rather, it suggests that the multi-partite entanglement could be spread among non-local correlations, rendering the operator used to access this information insufficient.
@@ IX Conclusions
While there are no perfectly clean samples in nature, none are infinitely disordered either. In this work, we set out to understand how one can experimentally access the affects of disorder on the entanglement of ground states within condensed matter systems. We utilized the quantum spin chain as a representative example to investigate the impact of bond disorder. In this setting, we focussed on distinguishing between Tomonaga-Luttinger liquid (TLL) and random singlet (RS) states through experimentally accessible entanglement measures, like concurrence, residual-tangle (RT) and quantum fisher information (QFI).
@@
We saw that both the TLL and RS states are multi-partite entangled. For the TLL, the observed multi-partite entanglement is a consequence of the uninhibited growth of correlations. On the other hand, such correlations are completely suppressed for the RS state. Consequently, the observation of multi-partite entanglement in the RS state is consequence of the localization of multi-partite clusters upto a crossover lengthscale. Above this regime, the relevant entanglement might indeed be 2-partite, consistent within the framework of strong disorder renormalization group (SDRG). To better characterize the effect of disorder, we introduced the inequality Eq. (60) which we found is also in agreement for the larger system sizes that were accessed in [90, 97].
@@
For quenched disorder, the order of average matters for convex roof measures. In particular, the distribution of pairwise entanglement, as characterized by concurrence, depends on how one performs the average. For the RS state, this distribution can transform from being shared between distant sites Fig. 5a to only among nearest neighbors Fig. 5b, as is the case in TLL. Since in INS one would access disorder-averaged correlations, it is the latter that one would observe rather than the former.
@@
For the TLL, it is well established that it can be described by a conformal field theory. We leveraged this connection to extract the associated central charge from the low-temperature behavior of concurrence. Conversely, for the RS state, we extracted the effective central charge from the low-temperature behavior of concurrence by employing disorder averaging accessible in experiments. This relies on appreciating that the order of disorder averaging alters the distribution of pairwise entanglement while preserving the universal information encoded within it. However, the existence of an effective central charge for the RS state remains conjectural, based on the assumption of a generalized c-theorem [92].
@@
From the experimental point of view, a more robust characteristic is to examine the low-temperature behavior of multi-partite entanglement witnesses like quantum Fisher information (QFI) or the equal time structure factor. We made use of the equal time structure-factor to show a marked difference in the rate of growth of multi-partite entanglement within the two states Fig. 7a. This scaling provides a clear distinction between the TLL and RS states, which can be observed experimentally.
@@
In the end, we would like to conclude, by pointing out that, in general some caution should be maintained especially when ruling out disorder-induced states using measures based on witnesses. However, when used carefully, these can provide important information about quantum spin systems in the presence of disorder.

### v3

@@ VIII.4 What can be useful experimentally?
For 1d critical systems, the low-temperature behavior of concurrence can access central charge information. In particular, the nearest neighbor concurrence of an $SU(2)$ symmetric phase can be related to the internal energy, and this allowed us to extract the central charge of the TLL. On the other hand, for RS phase, while the order of disorder-average effects how concurrence is distributed among the sites, the universal information about the phase should still be accessible if it is contained within the 2-partite entanglement. This universal information is the effective central charge. Since experimentally, the concurrence distribution for the RS phase can be nearest-neighbored just as the TLL, we thought that perhaps one might be able to access the effective central charge for the RS phase. This is based on the assumption that there exists a form of generalized c-theorem for the RS phase [98]. This is what we put to test in Fig. 6 and see some positive indications. However, we admit that this is conjectural at best.
@@
On the other hand, multi-partite entanglement witnesses like the QFI or the equal time structure factor are invaluable markers for not only identifying the existence of multi-partite entanglement but also how disorder affects their growth. We saw this by the marked difference in the low-temperature behavior of the equal time structure factor in the TLL and RS phase. This behavior can be experimentally useful in identifying the role of disorder in spin chains.
VIII.5 Limitations to detection of multi-partite entanglement
@@
Measures like QFI, when accessed via the dynamic structure factor, can only detect multi-partite entanglement if the underlying phase shows correlations peaked around specific $\mathbf{q}$ points. Due to the sum rule, this structure constraints the visibility of multi-partite entanglement detection at other $\mathbf{q}$ points. However, such peaked structure factors are not universal. For instance, in Kitaev and Kagome QSLs, there are no special $\mathbf{q}$ points with sharp features [105, 106, 107]. Therefore, multi-partite entanglement cannot be accessed using INS in these cases. This does not imply the absence of multi-partite entanglement in these phases, rather, it suggests that the multi-partite entanglement could be spread among non-local correlations, rendering the operator used to access this information insufficient.
@@ IX Conclusions
While there are no perfectly clean samples in nature, none are infinitely disordered either. In this work, we set out to understand how one can experimentally access the effects of disorder on the entanglement of ground states within condensed matter systems. We utilized the quantum spin chain as a representative example to investigate the impact of bond disorder. In this setting, we focussed on distinguishing between Tomonaga-Luttinger liquid (TLL) and random singlet (RS) phases through experimentally accessible entanglement measures, like concurrence, residual-tangle (RT) and quantum fisher information (QFI).
@@
We saw that both the TLL and RS phases are multi-partite entangled. For the TLL, the observed multi-partite entanglement is a consequence of the uninhibited growth of correlations. On the other hand, such correlations are completely suppressed for the RS phase. Consequently, the observation of multi-partite entanglement in the RS phase is consequence of the localization of multi-partite clusters up to a crossover length-scale. Above this regime, the relevant entanglement might indeed be 2-partite, consistent within the framework of strong disorder renormalization group (SDRG). To better characterize the effect of disorder, we introduced the inequality Eq. (60) which we found is also in agreement for the larger system sizes that were accessed in [94, 103].
@@
For quenched disorder, the order of average matters for convex roof measures. In particular, the distribution of pairwise entanglement, as characterized by concurrence, depends on how one performs the average. For the RS phase, this distribution can transform from being shared between distant sites Fig. 5a to only among nearest neighbors Fig. 5b, as is the case in TLL. Since in INS one would access disorder-averaged correlations, it is the latter that one would observe rather than the former.
@@
For the TLL, it is well established that it can be described by a conformal field theory. We leveraged this connection to extract the associated central charge from the low-temperature behavior of concurrence. Conversely, for the RS phase, we extracted the effective central charge from the low-temperature behavior of concurrence by employing disorder averaging accessible in experiments. This relies on appreciating that the order of disorder averaging alters the distribution of pairwise entanglement while preserving the universal information encoded within it. However, the existence of an effective central charge for the RS phase remains conjectural, based on the assumption of a generalized c-theorem [98].
@@
From the experimental point of view, a more robust characteristic is to examine the low-temperature behavior of multi-partite entanglement witnesses like quantum Fisher information (QFI) or the equal time structure factor. We made use of the equal time structure-factor to show a marked difference in the rate of growth of multi-partite entanglement within the two phases Fig. 7a. This scaling provides a clear distinction between the TLL and RS phases, which can be observed experimentally.
@@
In the end, we would like to conclude, by pointing out that, in general some caution should be maintained especially when ruling out disorder-induced phases using measures based on witnesses. However, when used carefully, these can provide important information about quantum spin systems in the presence of disorder.
@@ Acknowledgements
Appendix A Concurrence within pure and mixed states
In this Appendix we consider two different ways of evaluating concurrence for a pair of qubits, one appropriate for a mixed state, and one for pure state, with the aim of highlighting the difference between the two.
For a mixed state describing a pair of spins (qubits), concurrence is given by [67, 68],
$C_{ij}^{\sf mix}\equiv C(\rho^{\sf mix}_{ij})=\max\{0,\lambda_{1}-\lambda_{2}-\lambda_{3}-\lambda_{4}\}\;.$ (65)
where $\lambda_{\nu}$ are the solutions of
$\rho_{ij}\tilde{\rho}_{ij}\ket{v_{\lambda_{\nu}}}$ $=$ $\lambda_{\nu}^{2}\ket{v_{\lambda_{\nu}}}\;,$ (66)
arranged in descending order, with
$\tilde{\rho}_{ij}$ $=$ $(\sigma^{y}\otimes\sigma^{y})\rho^{*}_{ij}(\sigma^{y}\otimes\sigma^{y})\;,$ (67)
where reflects complex conjugation in a computational basis defined in terms of eigenstates of $\sigma^{z}$ (cf. Eq. (10), Eq. (11) and Eq. (12) of Section II).
In contrast, if the (reduced) density matrix describing the pair of spins reflects a pure state, then one can evaluate concurrence as [67, 68]
$C_{ij}^{\sf pure}=|\bra{\psi^{*}}\sigma_{i}^{y}\otimes\sigma_{j}^{y}\ket{\psi}|\;.$ | . (68)
To see that these two definitions lead to different answers when applied to many–body states with finite multipartite entanglement, we consider a 4–site Heisenberg spin chain with antiferromagnetic interactions ($J>0$), under periodic boundary conditions
${\mathcal{H}}=J({\bf S}_{1}\cdot{\bf S}_{2}+{\bf S}_{2}\cdot{\bf S}_{3}+{\bf S}_{3}\cdot{\bf S}_{4}+{\bf S}_{4}\cdot{\bf S}_{1})\;.$ (69)
The ground state is of this model is
|$\ket{\Psi}=\frac{1}{\sqrt{3}}\bigg{(}[12]\otimes[34]+[23]\otimes[41]\bigg{)}$ (70)
where
$[ij]=\frac{1}{\sqrt{2}}(\ket{0_{i}1_{j}}-\ket{1_{i}0_{j}})\;,$ (71)
is a singlet shared between sites $i$ and $j$.
We consider first the pair of sites $1$ and $2$. Tracing out the spins associated with sites $3$ and $4$, we find a reduced density matrix describing a mixed state
$\rho_{12}=\begin{pmatrix}1/12&0&0&0\\ 0&5/12&-1/3&0\\ 0&-1/3&5/12&0\\ 0&0&0&1/12\end{pmatrix}\;.$ (72)
The direct application of the result for a mixed state, Eq. (65), yields a concurrence
$C_{12}^{\sf mix}=\frac{1}{2}\;,$ (73)
reflecting the fact that a singlet, with concurrence $C=1$, and an unentangled state, with concurrence $C=0$, occur with equal probability for this pair of spins in Eq. (71).
If, instead, one were to blindly apply the result for pure state, Eq. (68), this would lead to an overestimate of the concurrence between the two spins
$C_{12}^{\sf pure}=\frac{2}{3}\;.$ (74)
Similarly, for the pair of sites $1$ and $3$, one should find
$C_{13}^{\sf mix}=0\;,$ (75)
since the wave function, Eq. (71), does not support any Bell state between this pair of spins. In contrast, if we apply the result for a pure state, Eq. (68), we find
$C_{13}^{\sf pure}=\frac{1}{3}\;,$ (76)
falsely diagnosing entanglement between this pair of spins.
The message, so far as disordered spin chains are concerned, is very clear: random singlet phases involve states with finite multipartite entanglement (cf. Section VI.2). For this reason, it is necessary to work the definition appropriate to a mixed state, Eq. (65), when calculating the concurrence associated with any given pair of spins.